OpenStudy (anonymous):

Simplify the following by first finding the lowest common denominator. 2/(x+3) + 3/(x-5)

OpenStudy (e.mccormick):

\[\frac{2}{x+3}+\frac{3}{x-5}\]LCD works the same in algebra as it does in fractions. For example: \[\frac{1}{2}+\frac{2}{3}=\frac{1\times 3}{2\times 3}+\frac{2\times 2}{3\times 2}\]

OpenStudy (e.mccormick):

If you do that same basic thing and multiply the (x-5) on the right to the top and bottom of the one on the left that will give it a denominator that can be added. At the same time, you mutliply (x+3) on the left times both the top and bottom of the right. Once you have that common denominatior, you can do the addition. Then it becomes a simplification problem.

OpenStudy (anonymous):

that means it simplifies to \[\frac{ 5x-1 }{ (x+3)(x-5)}\]

OpenStudy (e.mccormick):

Yah, and there are no factors below that match the above so that is a good answer. It is the fully factored form, which most teachers like.

OpenStudy (anonymous):

okay thankyou

OpenStudy (e.mccormick):

np. have fun!